A northbound train and a southbound train meet each other on parallel tracks heading in opposite directions. The northbound train travels 6 miles per hour faster than the southbound train. After 2 hours, they are 232 miles apart. At what speeds are the two trains traveling?

Answer by ikleyn(52835)

.

Let "r" be the rate of the slower train, in miles per hour.


Then the rate of the faster train is (r+6) miles per hour.


The total distance equation is


    2r + 2*(r+6) = 232   miles


saying that partial traveled distances sum up to the total separation distance.


The setup is just done, and what follows is the solution to the equation, step by step


    2r + 2r + 6 = 232

    4r          = 232 - 12

    4r          = 220

     r          = 220/4 = 55 miles per hour.


ANSWER.  The slower (the southbound) train's rate is 55 mph.

         The faster (the northbound) train's rate is 55 + 6 = 61 mph.

Solved.

--------------

For simple Travel @ Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.


-------------

Do you like my prose ?

I wrote my solution specially in this way to show that even the solution to a standard trivial routine problem
can be presented in an attractive form,  as a good tale story for children . . .


My other secret goal was to make you familiar with the terminology in this class of problems,
in order for you will be able to re-tell the story (the solution) to others.


And also because I love the Travel and Distance problems by very special love . . .